OpenStudy (anonymous):

The elevation of city A is 72 feet above sea level; the elevation of city C is 24 feet above sea level. The elevation of city D is 1 over 4 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. (5 points) Part B: Solve the expression and explain why the answer is negative or positive. (5 points)

OpenStudy (nisha_priya05):

elevation of D = 1/4 C= (1/4)* 24 =6 positive ans since elevation cannot be negative